# Towards a General-Purpose Zero-Shot Synthetic Low-Light Image and Video Pipeline

### v1

@@ Abstract
Low-light conditions pose significant challenges for both human and machine annotation. This in turn has led to a lack of research into machine understanding for low-light images and (in particular) videos. A common approach is to apply annotations obtained from high quality datasets to synthetically created low light versions. In addition, these approaches are often limited through the use of unrealistic noise models. In this paper, we propose a new Degradation Estimation Network (DEN), which synthetically generates realistic standard RGB (sRGB) noise without the requirement for camera metadata. This is achieved by estimating the parameters of physics-informed noise distributions, trained in a self-supervised manner. This zero-shot approach allows our method to generate synthetic noisy content with a diverse range of realistic noise characteristics, unlike other methods which focus on recreating the noise characteristics of the training data. We evaluate our proposed synthetic pipeline using various methods trained on its synthetic data for typical low-light tasks including synthetic noise replication, video enhancement, and object detection, showing improvements of up to 24% KLD, 21% LPIPS, and 62% AP50-95, respectively.
Index Terms:
@@
Synthetic Data, Low-Light, General-Purpose, Zero-Shot, Self-Supervised
1 Introduction
The performance of computer vision methods for various image and video processing tasks (e.g. classification, action recognition, object detection, object tracking, instance segmentation, etc.) has advanced significantly in recent years. This is due primarily to advances in deep learning methods and increased computational power, allowing deeper and more complex networks to be trained. The demand for video-based research is evidenced by the numerous international challenges in recent years  [1, 2, 3, 4, 5, 6, 7]. This has, in turn, enabled many practical applications including autonomous driving, digital surveillance, augmented/virtual reality (AR/VR), media production, and entertainment.
@@ 1 Introduction
Figure 1: Comparison of our proposed method with existing synthetic noise pipelines, where our model outperforms these methods across numerous benchmarks. Results are taken from Table I and normalized for visualization. For metrics where lower is better, indicated by $\downarrow$, the results are inverted.
@@
Despite these dramatic advances, research on processing videos captured under low-light conditions remains underdeveloped. Low light acquisition introduces numerous degradations such as low-contrast and motion blur, alongside sensor noise. All of these impact the performance of algorithms trained on normal-light data [8]. A further challenge is the lack of paired data, which is difficult to annotate accurately under low-light conditions without clear ground truth. As a result, only a limited number of datasets are available that offer adequate annotations for low-light video perception tasks [9, 10, 11, 12].
@@
To overcome this, a common approach is to first apply low-light video enhancement (LLVE) methods [13, 14, 15, 16] before performing the desired downstream task. This enables models trained on normal-light conditions to process the video without having been trained specifically on low-light. Low-light enhancement is, however, an ill-posed problem, where a single low-light input could be mapped to multiple potential outputs. This may introduce greater variability into the system, making it more difficult to solve a downstream task with a single solution. Moreover, adding extra pre-processing steps increases computational complexity, especially for deep-learning-based methods, thereby limiting real-time applications. In addition, imperfect pre-processing of input data may produce artifacts that deteriorate the performance of downstream models  [17, 18].
@@
An alternative approach is to generate synthetic data for training and evaluating the models. Mimicking real low-light conditions in this way removes the need to source suitable low-light datasets with the appropriate annotations; an existing clean dataset with annotations can be used instead. Although there are numerous pipelines available for creating synthetic low-light images and videos, each comes with limitations. Simpler pipelines often employ physics-informed distributions, which require manual selection of appropriate noise parameters  [19, 20] and fail to replicate sRGB-specific noise characteristics arising from in-camera processing pipelines. More complex pipelines may be tailored to specific types of noise, which significantly restricts their applications, or they may require camera-specific information that is not available in most publicly accessible datasets for downstream tasks.
In this context, we introduce a novel, general-purpose, zero-shot framework to address the limitations present in existing synthetic data pipelines. Our approach generates synthetic data with diverse, realistic noise characteristics, suitable for training models on new, noise-specific tasks. Crucially, our model operates without requiring training on individual datasets as it functions purely through inference on reference noisy data. This capability stems from our method of estimating a vector of noise parameters that maps to the corresponding physics-informed statistical distributions. By sampling from these distributions, our method can accurately replicate the noise characteristics found in any given reference input onto clean desired image or video data, enhancing the applicability and flexibility of the training data for noise-specific task models.
The main contributions are summarized as follows:
@@
  • Unlike the only existing zero-shot method proposed by [21], which is limited to Gaussian-distributed noise, our method offers a broad range of noise types and parameters, making it suitable for real-world applications.
@@
The proposed method has been evaluated across various tasks, including synthetic noise replication, video enhancement, and object detection, demonstrating improvements of up to 24% KLD, 21% LPIPS, and 62% AP50-95, respectively. Comparative analysis against established noise models, including Additive White Gaussian Noise (AWGN), Poisson-Gaussian noise (P-G) [19], and the Starlight noise generator [22], reveals our method’s superior performance, as shown in Fig. 1.
2 Related Work
2.1 RAW Noise Pipelines
@@
RAW image data inherently contains various types of noise, often modeled using statistical distributions that reflect the physical processes of image capture. Many researchers make use of these physical phenomena to synthesize noisy data for training and analysis. The most commonly modeled noise type is read noise, which is typically modeled using a Gaussian distribution. Other forms of noise, such as shot noise, banding noise, or quantization noise, are also represented using appropriate statistical models. While these methods are capable of modeling many noise types using only statistics, they require manual parameter selection to ensure a realistic output. This is evidenced in work by Wei et al. [23], where meticulous experimentation was required to estimate the noise parameters for five different cameras.
@@
More recently, there has been an increase in the use of deep-learning methods to synthesize realistic low-light images, without the need for parameter estimation [24, 25]. These methods are often ‘physics-guided’, allowing neural networks to learn the sampling of the relevant values for generating realistic low-light imagery. Cao et al. [24] developed a method that models a variety of different noise types, but it requires the availability of ISO information for the image. Meanwhile, Zhang et al. [25] proposed a more general-purpose pipeline, but solely focus on read and shot noise. Monakhova et al. [22] proposed a generative adversarial network (GAN) for synthesizing noisy RAW videos to train their denoiser. Similar to other pipelines, their method employs physics-guided noise, where the GAN initially samples from several statistical distributions with learned parameters for their dataset, followed by processing the video through a U-Net [26] to learn other complex noise types. However, their method requires re-training when applied to different cameras or datasets, highlighting a common limitation in the generalizability of current noise synthesis approaches.
2.2 sRGB Noise Pipelines
@@
Many synthetic pipelines predominantly use the RAW format to model noise based on physics principles. However, sRGB images inherently exhibit additional degradation types which need to be considered. One of the main degradation types is spatially-correlated noise, which is often caused by compression algorithms. Due to the complex nature of noise types like this, limited research has been conducted for synthetic noisy sRGB pipelines. Deep learning approaches have been explored, but these usually require further information about the image, such as camera parameters and ISO settings. For example, Kousha et al. [27] synthesized noisy sRGB images by leveraging normalizing flows, a family of generative models which use invertible transformations to translate data from one domain to another. Their method includes a specific linear flow layer conditioned on the camera metadata and gain (ISO) to ensure accurate synthetic noisy images.
@@
Fu et al. [28] introduced an alternative approach for generating noisy sRGB images that accounts for spatially-correlated noise without the need for ISO data of the captured image. They proposed a ‘gain estimation network’ that enables the model to autonomously synthesize noise without additional camera information. However, their method is still limited in terms of applications, as it is unable to generalize across different camera models, requiring re-training for each distinct camera.
2.3 Low-Light sRGB Pipelines
@@
Existing pipelines for generating low-light images and videos often rely on simpler models that primarily capture common noise types such as read and shot noise, and primarily focus on modifying the brightness and contrast of the data. This approach generally lacks the sophistication required to replicate realistic noise conditions effectively.
@@
Lv et al. [29] proposed a method for generating low-light images which mimics the image signal processing (ISP) pipeline in low-light photography. They first reduce the brightness of the images using a combination of linear and power transformations, before converting the images to a Bayer format. This synthetically creates a ‘RAW’ low-light image, where physics-informed noise types (i.e. read and shot noise) are applied, before converting back into sRGB format via demosaicing techniques. Cui et al. [30] also proposed a similar approach to simulating the ISP pipeline, but factor in further image transformations such as white balance and gamma correction to enhance realism.
@@
For videos, Zhou et al. [31] proposed a synthetic low-light pipeline focusing on motion blur due to low-light conditions. Unlike the simpler brightness adjustment techniques, they use a modified version of Zero-DCE [13] to reduce the brightness of the image, instead of simple linear or power transformations. However, only Gaussian and Poisson noise distributions are present in their noise model.
3 Proposed Method
Figure 2: The outline for training with our proposed Degradation Estimation Network (DEN), including its architecture. The architecture of the DEN consists of a U-Net [26] combined with a Multi-Layer Perceptron (MLP) head after the encoder to estimate the noise vector $\hat{\mathbf{v}}$ with denoising guidance.
Figure 3: The process for inferencing the DEN. A real noisy video is passed into the model to obtain the noise vector $\hat{\mathbf{v}}$, which accurately describes the noise characteristics in the video. The vector is used to sample from the relevant distributions onto the clean inputs to synthetically generate low-light videos.
@@ 3 Proposed Method
$\mathbf{x}_{d}=\alpha\mathbf{x}_{c}^{1/\gamma},$ (1)
@@
$\begin{split}\mathbf{x}_{n}&=\mathbf{x}_{d}+f(\mathbf{\hat{v}}),\\ \mathbf{\hat{v}}&=\text{DEN}(\mathbf{y}_{n}),\end{split}$ (2)
@@
3.1 Physics-Based Noise
@@
Read noise and shot noise are typical in low-light imagery  [32, 33]. Read noise arises from electrical noise inherent in the signal readout process from the camera sensor. It is a combination of a variety of noise types that can generally be modeled using a Gaussian distribution [33]. Shot noise, on the other hand, results from the quantum nature of light. During image capture, some photons may not have reached the sensor, leading to image degradation. This type of noise is particularly more noticeable in low-light scenes where photon counts are low. This effect can be modeled using a Poisson distribution. To simulate both types of noise, we employ the heteroscedastic noise model proposed by Foi et al. [19], which approximates the Poisson distribution as zero-mean Gaussian. The heteroscedastic noise model is beneficial as it can approximate shot noise, a type of signal-dependent noise that is not inherently additive, to be treated in an additive manner. The heteroscedastic noise model is mathematically described for a video frame of height $H$, width $W$ and channels $C$ as
$\mathbf{n}_{h}\sim\mathcal{N}(0,\sigma_{s}^{2}\mathbf{x}+\sigma_{r}^{2}),\quad\mathbf{n}_{h}\in\mathbb{R}^{C\times H\times W},$ (3)
@@
$\mathbf{n}_{q}\sim\mathcal{U}(0,\lambda_{q}),\quad\mathbf{n}_{q}\in\mathbb{R}^{C\times H\times W},$ (4)
@@
Banding noise refers to the horizontal or vertical lines that become prominent at high ISO, commonly used in low-light photography and videography to capture content without the need for long shutter speeds (which causes motion blur). This type of noise is highly camera-specific. Banding noise, like vertical banding as outlined by Monakhova et al. [22], is modeled using a Gaussian distribution centered at zero, with the variability of band positions encapsulated by the standard deviation $\sigma_{b}$. The noise model is represented as
$\mathbf{n}_{b}\sim\mathcal{N}(0,\sigma_{b}^{2}),\quad\mathbf{n}_{b}\in\mathbb{R}^{C\times 1\times W},$ (5)
where $\mathbf{n}_{b}$ is the banding noise map. Additionally, we include temporal banding noise $\mathbf{n}_{b_{t}}$ (as described in [22]), where banding noise maintains consistency throughout the video, characterized by a standard deviation of $\sigma_{b_{t}}$ to model noise persistence over time.
Periodic noise appears as a repeating pattern, often caused by electrical interference during the image capturing process. It is characterized by regularly spaced artifacts such as stripes or grids. Periodic noise is distinct from banding noise due to its strictly periodic nature. Unlike random noise, periodic noise can be effectively identified in the frequency domain, where it exhibits as distinct peaks. We follow the periodic noise implementation described in [22], which is formulated by
$\mathbf{n}_{p}[r,c]=\begin{cases}\mathcal{F}^{-1}(\mathbf{n}_{p_{1}}),&\text{if}\>\>c=0,\\ \mathcal{F}^{-1}(\mathbf{n}_{p_{2}}+\mathbf{n}_{p_{3}}i),&\text{if}\>\>c=W/4,\\ \mathcal{F}^{-1}(\mathbf{n}_{p_{2}}-\mathbf{n}_{p_{3}}i),&\text{if}\>\>c=3W/4,\\ 0,&\text{otherwise},\end{cases}$ (6)
@@
3.2 Degradation Estimation Network (DEN)

### v2

@@ Abstract.
Low-light conditions pose significant challenges for both human and machine annotation. This in turn has led to a lack of research into machine understanding for low-light images and (in particular) videos. A common approach is to synthesize low-light datasets from existing datasets with high-quality annotations. However, these approaches are often limited through the use of unrealistic noise models, or require metadata information often unavailable in public datasets. In this paper, we propose a new Degradation Estimation Network (DEN), which synthetically generates realistic standard RGB (sRGB) noise without the requirement for camera metadata. This is achieved by estimating the parameters of physics-informed noise distributions, trained in a self-supervised manner. This zero-shot approach allows our method to generate synthetic noisy content with a diverse range of realistic noise characteristics, unlike other methods which focus on recreating the noise characteristics of the training data. We evaluate our proposed synthetic pipeline using various methods trained on its synthetic data for typical low-light tasks including synthetic noise replication, video enhancement, and object detection, showing improvements of up to 24% KLD, 21% LPIPS, and 62% AP50-95, respectively.
@@
Synthetic data, Low-Light, Degradations, General-Purpose, Zero-Shot, Self-Supervised
journalyear: 2025copyright: ccconference: Proceedings of the 3rd International Workshop on Multimedia Content Generation and Evaluation: New Methods and Practice; October 27–28, 2025; Dublin, Irelandbooktitle: Proceedings of the 3rd International Workshop on Multimedia Content Generation and Evaluation: New Methods and Practice (McGE ’25), October 27–28, 2025, Dublin, Irelanddoi: 10.1145/3746278.3759376isbn: 979-8-4007-2060-4/2025/10ccs: Computing methodologies Image and video acquisitionccs: Computing methodologies Computational photographyccs: Computing methodologies Image processingccs: Computing methodologies Scene understandingccs: Computing methodologies Reconstructionccs: Computing methodologies Object detectionccs: Computing methodologies Tracking
1. Introduction
@@
Figure 1. Comparison of our proposed method with existing synthetic noise pipelines, where our model outperforms these methods across numerous benchmarks. Results are taken from Table 1 and normalized for visualization. For metrics where lower is better, indicated by $\downarrow$, the results are inverted.
[…]
Radar plot using several metrics where our model outperforms the synthetic noise pipelines.
The performance of computer vision methods for various image and video processing tasks (e.g. classification, action recognition, object detection, object tracking, instance segmentation, etc.) has advanced significantly in recent years. This is due primarily to advances in deep learning methods and increased computational power, allowing deeper and more complex networks to be trained. The demand for video-based research is evidenced by the numerous international challenges in recent years  (Sigurdsson et al., 2016; Perazzi et al., 2016; Xu et al., 2018; Damen et al., 2018; Yang et al., 2019; Huang et al., 2021; Anantrasirichai et al., 2025). This has, in turn, enabled many practical applications including autonomous driving, digital surveillance, augmented/virtual reality (AR/VR), media production, and entertainment.
@@
Despite these dramatic advances, research on processing videos captured under low-light conditions remains underdeveloped. Low-light acquisition introduces numerous degradations such as low-contrast and motion blur, alongside sensor noise. All of these impact the performance of algorithms trained on normal-light data (Yi and Anantrasirichai, 2024). A further challenge is the lack of paired data, which is difficult to annotate accurately under low-light conditions without clear ground truth. As a result, only a limited number of datasets are available that offer adequate annotations for low-light video perception tasks (Ye et al., 2022; Li et al., 2024; Wang et al., 2024; Liu et al., 2024).
@@
To overcome this, a common approach is to first apply low-light video enhancement (LLVE) methods (Guo et al., 2020; Wang et al., 2021; Fu et al., 2023b; Lin et al., 2024b) before performing the desired downstream task. This enables models trained on normal-light conditions to process the video without having been trained specifically on low-light. Low-light enhancement is, however, an ill-posed problem, where a single low-light input could be mapped to multiple potential outputs. This may introduce greater variability into the system, making it more difficult to solve a downstream task with a single solution. Moreover, adding extra pre-processing steps increases computational complexity, especially for deep-learning-based methods, thereby limiting real-time applications. In addition, imperfect pre-processing of input data may produce artifacts that deteriorate the performance of downstream models  (Lin et al., 2025; Chen and Ma, 2025).
@@
An alternative approach is to generate synthetic data for training and evaluating the models. Using synthetic pipelines removes the need to source suitable low-light datasets with the appropriate annotations; an existing clean dataset with annotations can be used instead. Although there are numerous pipelines available for creating synthetic low-light images and videos, each comes with limitations. Traditional pipelines often employ physics-informed distributions, which require manual selection of appropriate noise parameters (Foi et al., 2008; Anantrasirichai et al., 2015) and fail to replicate sRGB-specific noise characteristics arising from in-camera processing pipelines while more complex pipelines may be tailored to noise types from specific datasets, which significantly restricts their applications, or they may require camera-specific information that is not available in most publicly accessible datasets for downstream tasks.
In this context, we introduce a novel, general-purpose, zero-shot framework to address the limitations present in existing synthetic data pipelines. Our approach is capable of capturing a diverse range of realistic degradation types, by estimating a vector of noise parameters which is used to sample noise maps from the corresponding physics-informed distributions. The main contributions are summarized as follows:
@@
  • Unlike the only existing zero-shot deep-learning method proposed by (Luo et al., 2023), which is limited to Gaussian-distributed noise, our method offers a broad range of noise types and parameters, making it suitable for real-world applications.
@@
The proposed method has been evaluated across various tasks, including synthetic noise replication, video enhancement, and object detection, demonstrating improvements of up to 24% KLD, 21% LPIPS, and 62% AP50-95, respectively. Comparative analysis against established noise models, including Additive White Gaussian Noise (AWGN), Poisson-Gaussian noise (P-G) (Foi et al., 2008), and the Starlight noise generator (Monakhova et al., 2022), reveals our method’s superior performance, as shown in Fig. 1.
2. Related Work
2.1. RAW Noise Pipelines
@@
RAW image data inherently contains various types of noise, often modeled using statistical distributions that reflect the physical processes of image capture. Many researchers make use of these physical phenomena to synthesize noisy data for training and analysis. The most commonly modeled noise type is read noise, which is typically modeled using a Gaussian distribution. Other forms of noise, such as shot noise, banding noise, or quantization noise, are also represented using appropriate statistical models. While these methods are capable of modeling many noise types using only statistics, they require manual parameter selection to ensure a realistic output. This is evidenced in work by Wei et al. (Wei et al., 2021), where meticulous experimentation was required to estimate the noise parameters for five different cameras.
@@
More recently, there has been an increase in the use of deep-learning methods to synthesize realistic low-light images, without the need for parameter estimation (Cao et al., 2023; Zhang et al., 2023). These methods are often described as ‘physics-guided’, incorporating physics-informed distributions within the neural networks for generating realistic low-light imagery. Cao et al. (Cao et al., 2023) developed a method that models a variety of different noise types, but it requires the availability of ISO information for the image. Meanwhile, Zhang et al. (Zhang et al., 2023) proposed a more general-purpose pipeline, without the need for camera metadata, but their method is restricted to their training data. Monakhova et al. (Monakhova et al., 2022) proposed a generative adversarial network (GAN) for synthesizing noisy RAW videos to train their video denoiser. Similar to other physics-guided pipelines, their generator initially samples from several statistical distributions with learned parameters for their dataset, followed by processing the video through a U-Net (Ronneberger et al., 2015) to learn other complex noise types. However, like (Zhang et al., 2023), their method requires re-training when applied to different cameras or datasets, highlighting a common limitation in the generalizability of current noise synthesis approaches.
2.2. sRGB Noise Pipelines
@@
Many synthetic pipelines predominantly use the RAW format to model noise based on physics principles. However, sRGB images inherently exhibit additional degradation types which need to be considered. One of the main noise types is spatially-correlated noise, which is often caused by compression algorithms, denoising algorithms and other processes found in the image signal processing (ISP) pipeline. Due to the complex nature of noise types like this, limited research has been conducted for synthetic noisy sRGB pipelines. Deep learning approaches have been explored, but these usually require further information about the image, such as camera parameters and ISO settings. For example, Kousha et al. (Kousha et al., 2022) synthesized noisy sRGB images by leveraging normalizing flows, a family of generative models which use invertible transformations to translate data from one domain to another. Their method includes a specific linear flow layer conditioned on the camera metadata and gain (ISO) to ensure accurate synthetic noisy images.
@@
Fu et al. (Fu et al., 2023a) introduced an alternative approach for generating noisy sRGB images that accounts for spatially-correlated noise without the need for ISO data of the captured image. They proposed a ‘gain estimation network’ that enables the model to autonomously synthesize noise without additional camera information. However, their method is still limited in terms of applications, as it is unable to generalize across different camera models, requiring re-training for each distinct camera.
2.3. Low-Light sRGB Pipelines
@@
Existing pipelines for generating low-light images and videos often rely on simple noise models that primarily capture common noise types such as read and shot noise, and primarily focus on modifying the brightness and contrast of the data. This approach generally lacks the sophistication required to replicate realistic noise conditions effectively.
@@
Lv et al. (Lv et al., 2021) proposed a method for generating low-light images which mimics the ISP pipeline in low-light photography. They first reduce the brightness of the images using a combination of linear and power transformations, before converting the images to a Bayer format. This synthetically creates a ‘RAW’ low-light image, where physics-informed noise types (i.e. read and shot noise) are applied, before converting back into sRGB format via demosaicing techniques. Cui et al. (Cui et al., 2021) also proposed a similar approach to simulating the ISP pipeline, but factor in further image transformations such as white balance and gamma correction to enhance realism. Luo et al. (Luo et al., 2023) proposed a deep-learning based method for synthesizing low-light images, by training a network to estimate adjustment maps to reduce the brightness of the input images, conditioned by exposure map. However, their method only draws noise maps from Gaussian distributions, and does not consider signal-dependent noise.
@@
For videos, Zhou et al. (Zhou et al., 2022) proposed a synthetic low-light pipeline focusing on blur degradations due to low-light conditions. Unlike the simpler brightness adjustment techniques, they use a modified version of Zero-DCE (Guo et al., 2020) to reduce the brightness of the image, instead of simple linear or power transformations. They use a pre-existing method, CycleISP (Zamir et al., 2020), for synthesizing noise.
3. Proposed Method
Figure 2. Overview of the proposed framework. The Degradation Estimation Network (DEN) is based on a U-Net (Ronneberger et al., 2015) with an MLP head attached to the encoder, used to predict a noise vector $\hat{\mathbf{v}}$ with denoising guidance. In training, the illumination reducer and noise simulator function $f(\cdot)$ are used to create simulated noisy inputs for DEN. In inference, a real low-light input $\mathbf{y}_{n}$ is used to obtain a noise vector to be applied onto the clean input.
[…]
The pipeline of the proposed framework, showing the training and inference steps of using our Degradation Estimation Network (DEN) to predict the noise vector $\hat{\mathbf{v}}$ which is used to recreate the input noise characteristics.
@@ 3. Proposed Method
(1) $\mathbf{x}_{d}=\alpha\mathbf{x}_{c}^{1/\gamma},$
@@
(2) $\begin{split}\mathbf{x}_{n}&=\mathbf{x}_{d}+f(\mathbf{\hat{v}}),\\ \mathbf{\hat{v}}&=\text{DEN}(\mathbf{y}_{n}),\end{split}$
@@
3.1. Physics-Based Noise
@@
Heteroscedastic noise.
Read and shot noise are typical in low-light imagery  (El Gamal and Eltoukhy, 2005; Boncelet, 2009). Read noise arises from electrical noise inherent in the signal readout process from the camera sensor. It is a combination of a variety of noise types that can generally be modeled using a Gaussian distribution (Boncelet, 2009). Shot noise, on the other hand, results from the quantum nature of light. During image capture, some photons may not have reached the sensor, leading to image degradation. This type of noise is particularly more noticeable in low-light scenes where photon counts are low. This effect can be modeled using a Poisson distribution. To simulate both types of noise, we employ the heteroscedastic noise model proposed by Foi et al. (Foi et al., 2008), which approximates the Poisson distribution as zero-mean Gaussian. The heteroscedastic noise model is beneficial as it can approximate shot noise, a type of signal-dependent noise that is not inherently additive, to be treated in an additive manner. The heteroscedastic noise model is mathematically described for a video frame of height $H$, width $W$ and channels $C$ as
(3) $\mathbf{n}_{h}\sim\mathcal{N}(0,\sigma_{s}^{2}\mathbf{x}+\sigma_{r}^{2}),\quad\mathbf{n}_{h}\in\mathbb{R}^{C\times H\times W},$
@@
(4) $\mathbf{n}_{q}\sim\mathcal{U}(0,\lambda_{q}),\quad\mathbf{n}_{q}\in\mathbb{R}^{C\times H\times W},$
@@
Banding noise.
Banding noise refers to the horizontal or vertical lines that become prominent at high ISO, commonly used in low-light photography and videography to capture content without the need for long shutter speeds (which causes motion blur). This type of noise is highly camera-specific. Banding noise, like vertical banding as outlined by Monakhova et al. (Monakhova et al., 2022), is modeled using a Gaussian distribution centered at zero, with the variability of band positions encapsulated by the standard deviation $\sigma_{b}$. The noise model is represented as
(5) $\mathbf{n}_{b}\sim\mathcal{N}(0,\sigma_{b}^{2}),\quad\mathbf{n}_{b}\in\mathbb{R}^{C\times 1\times W},$
where $\mathbf{n}_{b}$ is the banding noise map. Additionally, we include temporal banding noise $\mathbf{n}_{b_{t}}$ (as described in (Monakhova et al., 2022)), where banding noise maintains consistency throughout the video, characterized by a standard deviation of $\sigma_{b_{t}}$ to model noise persistence over time.
Periodic noise.
Periodic noise appears as a repeating pattern, often caused by electrical interference during the image capturing process. It is characterized by regularly spaced artifacts such as stripes or grids. Periodic noise is distinct from banding noise due to its strictly periodic nature. Unlike random noise, periodic noise can be effectively identified in the frequency domain, where it exhibits as distinct peaks. We follow the periodic noise implementation described in (Monakhova et al., 2022), which is formulated by
(6) $\mathbf{n}_{p}[r,c]=\begin{cases}\mathcal{F}^{-1}(\mathbf{n}_{p_{1}}),&\text{if}\>\>c=0,\\ \mathcal{F}^{-1}(\mathbf{n}_{p_{2}}+\mathbf{n}_{p_{3}}i),&\text{if}\>\>c=W/4,\\ \mathcal{F}^{-1}(\mathbf{n}_{p_{2}}-\mathbf{n}_{p_{3}}i),&\text{if}\>\>c=3W/4,\\ 0,&\text{otherwise},\end{cases}$
@@
3.2. Degradation Estimation Network (DEN)